# 22-quasi-perfect Lee codes and abelian Ramanujan graphs: a new construction and relationship

### v1

@@ Abstract.
In this paper, we obtain a new explicit family of $2$-quasi-perfect Lee codes of arbitrarily large length. Our construction is based on generating sets of abelian (almost) Ramanujan graphs obtained by Forey, Fresán, Kowalski and Wigderson. Also, we develop a relationship between certain abelian Ramanujan graphs and $2$-quasi-perfect Lee codes obtained by Mesnager, Tang and Qi.
@@ 1. Introduction
Error-correcting codes constitute the mathematical backbone of reliable communication and data storage. While the theory of codes equipped with the Hamming distance has reached a mature stage—characterized by deep connections to algebraic geometry and combinatorics—the study of codes under the Lee distance presents distinct theoretical challenges and opportunities. The Lee distance, which measures the modular distance between symbols, naturally captures the topology of phase-shift keying (PSK) modulation, circular signal constellations, and high-density data storage systems such as multilevel flash memories and DNA storage (e.g. [8, 10, 16]). Unlike the Hamming distance, which treats all non-zero errors equally, the Lee distance is sensitive to the magnitude of the error, making it inherently “arithmetic” rather than purely combinatorial. Historically, the search for optimal Lee codes has been driven by the quest for perfect Lee codes. The seminal work of Golomb and Welch [9] laid the foundation by formulating the sphere-packing bound for the Lee distance and conjecturing the non-existence of perfect Lee codes for packing radius $t\geq 2$ and length $n>2$. This “Golomb-Welch conjecture” remains a central open problem in combinatorial coding theory. From a broader theoretical perspective, however, the focus has shifted from the elusive perfect codes to the construction of “good” codes.
@@
As a class of such codes, quasi-perfect Lee codes have been studied in the literature. Constructing linear quasi-perfect Lee codes over represents a fascinating intersection of number theory, algebraic geometry, combinatorics, and coding theory. The fundamental mathematical challenge lies in the tension between the algebraic structure of the base field (used to define the code) and the geometric nature of the Lee distance. While linear codes are subspaces defined by algebraic operations, the Lee weight is not invariant under arbitrary field automorphisms. Consequently, constructing quasi-perfect Lee codes requires identifying specific algebraic structures such as algebraic curves that induce favorable weight distributions in the integer domain. Recent advances have demonstrated that techniques from algebraic geometry, particularly the estimation of character sums (exponential sums) and the number of rational points on curves, are powerful tools for this purpose. The celebrated Hasse-Weil bound (e.g. [17]), for instance, allows for precise estimates of the minimum distance and the covering radius of codes defined by polynomial images. In this context, Camarero and Martínez [3] constructed $2$-quasi-perfect $p$-ary Lee codes of length $2\lfloor\frac{p}{4}\rfloor$ for arbitrary primes $p\equiv\pm 5\pmod{12}$, which proves the existence of quasi-perfect Lee codes for arbitrarily large length. For an odd prime $p$, Mesnager, Tang, and Qi [15] formulated an equivalent condition for the existence of $2$-quasi-perfect $p$-ary linear codes of length $n$, and provided several explicit families of $2$-quasi-perfect Lee codes by generalizing the Camarero-Martínez construction, and analyzing the solution sets of specific polynomial equations. These works exemplify how deep arithmetic results can be translated into coding theoretic gains. Moreover, the Camarero-Martínez and Mesnager-Tang-Qi constructions also produce Cayley graphs over the additive groups of finite fields that are abelian Ramanujan graphs (as these groups are abel groups), which bridges between the theory of Lee codes and spectral graph theory.
@@
To the best of our knowledge, constructing quasi-perfect Lee codes of arbitrarily large length remains a formidable open problem. A unified theoretical framework that rationalizes why certain algebraic structures yield such codes is notably absent; consequently, existing constructions often appear as isolated families derived from specific polynomials without clear intuitive grounding. Furthermore, while the small diameter of abelian Ramanujan graphs inherently favors the covering radius of associated Lee codes, achieving perfect or quasi-perfect status necessitates stricter structural constraints. Determining precisely which abelian Ramanujan graphs yield these optimal codes remains a significant open question.
@@
In this paper, we first exhibit a new and simple construction for $2$-quasi-perfect linear Lee codes of arbitrarily large length. Specifically, we investigate the $p$-ary linear code derived from the set $H_{q}\coloneq\{(a,a^{3})\mid a\in\mathbb{F}_{q}^{*}\}$, where $p\geq 5$ is a prime and $q=p^{k}$, and prove that this code is a $2$-quasi-perfect code of length $\frac{q-1}{2}$ and dimension $\frac{q-1}{2}-2k$ provided that $q\geq 14$ (Theorem 4). It is worth mentioning that this construction comes from a recent construction of abelian almost Ramanujan graphs by Forey, Fresán, Kowalski and Wigderson [7], and the structure of the codes and associated Cayley graphs are different from known ones. Moreover, we bridge the Mesnager-Qi-Tang construction of Lee codes to well-studied abelian Ramanujan graphs in combinatorics and number theory, namely, Li’ s graphs introduced by Winnie Li, and finite Euclidean graphs. This provides a unified description of the families of Lee codes by the Mesnager-Qi-Tang construction (Propositions 16 and 18).
@@
The remainder of the paper is organized as follows. Section 2 reviews preliminaries on finite fields, Lee codes, Cayley and Ramanujan graphs. Section 4 details our construction and the proof of the $2$-quasi-perfectness employing algebraic geometry. Section 5 discusses a connection between the Mesnager-Qi-Tang codes and abelian Ramanujan graphs, namely, Li’s graphs and finite Euclidean graphs.
@@ 2.3. Cayley graphs and Ramanujan graphs
For a finite group $G$ with identity $1$, and an inverse-closed subset $S\subset G\setminus\{1\}$, the Cayley graph $Cay(G,S)$ on $G$ with respect to $S$ is the graph with vertex set $G$ in which two vertices $g$ and $h$ are adjacent if and only if $gh^{-1}\in S$. Note that $Cay(G,S)$ is an undirected $|S|$-regular graph.
The adjacency matrix $A_{X}$ of a simple $N$-vertex graph $X$ is an $N\times N$ $\{0,1\}$-matrix where the $(u,v)$-entry is $1$ if and only if $u$ and $v$ are adjacent in $X$, and hence, the matrix is symmetric, implying that all eigenvalues are real. By the Perron-Frobenius theorem, for any $d$-regular graph $X$, all eigenvalues of $A_{X}$ are in the interval [d,d][-d,d], where the largest eigenvalue is $d$, and d-d is an eigenvalue if and only if $X$ is bipartite. We call $\pm d$ trivial eigenvalues, and set $\lambda(X)$ as the largest absolute value of non-trivial eigenvalues of $A_{X}$. A $d$-regular NN-vertex graph $X$ is called a Ramanujan graph if $\lambda(X)\leq 2\sqrt{d-1}$. Similarly, we say $X$ is an almost Ramanujan graph if $\lambda(X)\leq(2+o(1))\sqrt{d-1}$ as $N\to\infty$ If $d$ is fixed, then $\liminf_{N\to\infty}\lambda(X)\geq 2\sqrt{d-1}$ by the Alon-Boppana bound (e.g. [5]), and even if $d$ grows as $N\to\infty$ (say, $\omega(1)=d<\frac{N}{2}$), one can prove that $\lambda(X)=\Omega(\sqrt{d})$ as $N\to\infty$. Hence, Ramanujan graphs realize the optimally small non-trivial eigenvalues in absolute value (up to constant). (Almost) Ramanujan graphs enjoy very rich properties such as high global connectivity, pseudo-randomness, and rapid mixing. An abelian Ramanujan graph is a Cayley graph over an abelian group which is a Ramanujan graph. For a more detailed exposition of Cayley and Ramanujan graphs, see, for instance, [5].
@@ 4. A new explicit construction of $2$-quasi-perfect Lee codes
For $p=19\equiv 7\pmod{12}$ and $k=1$, the parity check matrix of the code $\mathcal{C}(\mathbb{F}_{13}^{2},H_{13})$ is
@@
Note that $a=0$ if and only if y=xy=-x, equivalent to $b=0$, implying that there are exactly $\frac{q-1}{2}$ unordered pair $\{x,y\}$ satisfying (4.1) with $(a,b)=(0,0)$. If $a\neq 0$, then $(a,b)$ corresponds to an unordered pair $\{x,y\}$ satisfying (4.1) and $x+y\neq 0$ by a one-to-one mapping. Indeed, by (4.1), we have $xy=\frac{a^{3}-b}{3a}$, and hence $x,y$ must be roots of the quadratic equation $t^{2}-at+\frac{a^{3}-b}{3a}=0$. Now we obtain $H_{q}^{(2)}$ as the number of $\{x,y\}$ with $x+y\neq 0$ is
${}_{q-1}H_{2}-\frac{q-1}{2}=\binom{(q-1)+2-1}{2}-\frac{q-1}{2}=\frac{(q-1)^{2}}{2}.$
Here notice that $(0,0)\in H_{q}^{(2)}$, implying that $H_{q}^{(0)}\subset H_{q}^{(2)}$. Thus, the proof completes after showing that $H_{q}\cap H_{q}^{(2)}=\emptyset$ as $H_{q}^{(1)}=\#H_{q}=q-1$. Suppose that there exists $(a,a^{3})\in H_{q}$ such that there are $x,y\in\mathbb{F}_{q}^{*}$ with
@@
This implies that $3axy=0$, which is impossible as $x,y,a\neq 0$. ∎
@@
Suppose $b\neq a^{3}$ and $F_{a,b}$ is not absolutely irreducible. As the leading term of $F_{a,b}$ is $xy(x+y)$, $F_{a,b}$ must have a linear factor which is either x+ycx+y-c, xcx-c or ycy-c for some constant $c$. If x+ycx+y-c is a factor of $F_{a,b}$, then we have $(c-a)(xy+ac)+\frac{b-a^{3}}{3}=0$ which must hold for any $x$ and $y$. Hence it must hold that $c=a$ and $b=a^{3}$, which cannot occur by the assumption. If xcx-c is a factor of $F_{a,b}$, then for any $y$, $(y+c-a)((c+a)y+ac)+\frac{b-a^{3}}{3}=0$ holds. Hence we have $c+a=0$, (ca)(c+a)+ac=0(c-a)(c+a)+ac=0 and $(c-a)ac+\frac{b-a^{3}}{3}=0$, which implies $b=a^{3}$, contradiction. By the same discussion, ycy-c cannot be a factor of $F_{a,b}$, completing the proof. ∎
@@
If $b=a^{3}$, as shown in Lemma 12, the curve $C_{a,b}$ is the union of the line $x+y=a$ and the hyperbolic curve $h_{a}:xy+a(x+y)=0$. As $a\neq 0$ in this case, it suffices to consider the points $(x,y)$ of the curve $h_{a}$ with $x,y\neq 0$ and $x+y\neq a$. Any of the lines $x=0$ and $y=0$ intersects with $h_{a}$ at only $(0,0)$, and the line $x+y=a$ intersects with $h_{a}$ at no more than two points. As $h_{a}$ has totally q1q-1 points, there are at least q4q-4 desired points $(x,y)$.
@@ 2. $(2)$
$Q(x,y)=\left(x+2^{-1}y\right)^{2}+3\cdot 4^{-1}y^{2}.$
Let $u=x+2^{-1}y$ and $v=y$. Then $Q(x,y)$ is equivalent to the diagonal form $u^{2}+3\cdot 4^{-1}v^{2}$.
  • If 3-3 is a square in $\mathbb{F}_{q}$, then $3\cdot 4^{-1}$ can be written as $s^{2}$ for some $s\in\mathbb{F}_{q}^{*}$. Thus, $Q$ is equivalent to $u^{2}-s^{2}v^{2}=(u-sv)(u+sv)$, which is equivalent to $Q_{-}(x,y)=xy$. In this case, $\mathcal{C}(\mathbb{F}_{q}^{2},E_{Q})$ coincides with $\mathcal{C}(\mathbb{F}_{q}^{2},E_{Q_{-}})$ up to coordinate permutations and sign changes.
  • If 3-3 is a non-square in $\mathbb{F}_{q}$, then $3\cdot 4^{-1}$ is a non-square (since $4^{-1}$ is a square). Let $\delta=-3\cdot 4^{-1}$. Then $Q$ is equivalent to $u^{2}-\delta v^{2}$, which is exactly $Q_{+}(u,v)$. In this case, $\mathcal{C}(\mathbb{F}_{q}^{2},E_{Q})$ coincides with $\mathcal{C}(\mathbb{F}_{q}^{2},E_{Q_{+}})$ up to coordinate permutations and sign changes.

### v2

@@ Abstract.
This paper presents a new explicit infinite family of 2-quasi-perfect $p$-ary Lee codes of length $\frac{q-1}{2}$ and dimension $\frac{q-1}{2}-2k$ for $q=p^{k}\geq 14$, $p\geq 5$ a prime. Our codes are derived from the generating set $H_{q}=\{(a,a^{3})\mid a\in\mathbb{F}_{q}^{*}\}$ of the finite field $\mathbb{F}_{q^{2}}$. Furthermore, we bridge between 2-quasi-perfect Lee codes constructed by Mesnager, Tang, and Qi and well-known abelian Ramanujan graphs, specifically Li’s graphs and finite Euclidean graphs, providing a unified theoretical framework for these families.
@@ 1. Introduction
A Lee code is an error-correcting code under the Lee distance. While finding perfect Lee codes is a classical problem, from a broader theoretical perspective, the focus has shifted to the construction of “good” codes, such as quasi-perfect Lee codes. Constructing linear quasi-perfect Lee codes over finite fields represents a fascinating intersection of coding theory, number theory, algebraic geometry, and combinatorics. In this context, Camarero and Martínez [3] constructed $2$-quasi-perfect $p$-ary Lee codes of length $2\lfloor\frac{p}{4}\rfloor$ for arbitrary primes $p\equiv\pm 5\pmod{12}$, which proves the existence of quasi-perfect Lee codes for arbitrarily large length. For an odd prime $p$, Mesnager, Tang, and Qi [15] formulated an equivalent condition for the existence of $2$-quasi-perfect $p$-ary linear codes of length $n$, and provided several explicit families of $2$-quasi-perfect Lee codes by generalizing the Camarero-Martínez construction and analyzing the solution sets of specific polynomial equations. Moreover, the Camarero-Martínez and Mesnager-Tang-Qi constructions also produce Cayley graphs over the additive groups of finite fields that are abelian Ramanujan graphs. It is worth mentioning that these graphs are optimal pseudorandom graphs in a sense of spectral gap, where their short diameters crucially affect the covering property of the corresponding Lee codes.
@@
Nevertheless, to the best of our knowledge, constructing quasi-perfect Lee codes of arbitrarily large length remains a challenging open problem. A unified theoretical framework that rationalizes why certain algebraic structures yield such codes is notably absent; consequently, existing constructions often appear as isolated families derived from specific polynomials without clear intuitive grounding. Furthermore, while the small diameter of abelian Ramanujan graphs inherently favors the covering radius of associated Lee codes, achieving perfect or quasi-perfect status necessitates stricter structural constraints. Determining precisely which abelian Ramanujan graphs yield these optimal codes remains a significant open question.
@@
In this paper, we first exhibit a new and simple construction for $2$-quasi-perfect linear Lee codes of arbitrarily large length. Specifically, we investigate the $p$-ary linear code derived from the set $H_{q}\coloneq\{(a,a^{3})\mid a\in\mathbb{F}_{q}^{*}\}$, where $p\geq 5$ is a prime and $q=p^{k}$, and prove that this code is a $2$-quasi-perfect code of length $\frac{q-1}{2}$ and dimension $\frac{q-1}{2}-2k$ provided that $q\geq 14$ (Theorem 4). It is worth mentioning that this construction comes from a recent construction of abelian almost Ramanujan graphs by Forey, Fresán, Kowalski and Wigderson [7], and the structure of the codes and associated Cayley graphs are different from known ones. Moreover, we bridge the Mesnager-Qi-Tang construction of Lee codes to well-studied abelian Ramanujan graphs in combinatorics and number theory, namely, Li’ s graphs introduced by Winnie Li ([11]), and finite Euclidean graphs ([14]). This provides a unified description of the families of Lee codes by the Mesnager-Qi-Tang construction (Propositions 16 and 18).
@@
The remainder of the paper is organized as follows. Section 2 reviews preliminaries on finite fields, Lee codes, Cayley and Ramanujan graphs. Section 4 details our construction and the proof of the $2$-quasi-perfectness employing algebraic geometry. Section 5 discusses a connection between the Mesnager-Qi-Tang codes and abelian Ramanujan graphs, namely, Li’s graphs and finite Euclidean graphs. We conclude this paper with several remarks in Section 6.
@@ 2.3. Cayley graphs and Ramanujan graphs
For a finite group $G$ with identity $1$, and an inverse-closed subset $S\subset G\setminus\{1\}$ (which means that $s^{-1}\in S$ for any $s\in S$), the Cayley graph $Cay(G,S)$ on $G$ with respect to $S$ is the graph with vertex set $G$ in which two vertices $g$ and $h$ are adjacent if and only if $gh^{-1}\in S$. Note that $Cay(G,S)$ is an undirected $|S|$-regular graph.
The adjacency matrix $A_{X}$ of a simple $N$-vertex graph $X$ is an $N\times N$ $\{0,1\}$-matrix where the $(u,v)$-entry is $1$ if and only if $u$ and $v$ are adjacent in $X$, and hence, the matrix is symmetric, implying that all eigenvalues are real. By the Perron-Frobenius theorem, for any $d$-regular graph $X$, all eigenvalues of $A_{X}$ are in the interval [d,d][-d,d], where the largest eigenvalue is $d$, and d-d is an eigenvalue if and only if $X$ is bipartite. We call $\pm d$ trivial eigenvalues, and set $\lambda(X)$ as the largest absolute value of non-trivial eigenvalues of $A_{X}$. A $d$-regular NN-vertex graph $X$ is called a Ramanujan graph if $\lambda(X)\leq 2\sqrt{d-1}$. Similarly, we say $X$ is an almost Ramanujan graph if $\lambda(X)\leq(2+o(1))\sqrt{d-1}$ as $N\to\infty$ If $d$ is fixed, then it is well-known that $\liminf_{N\to\infty}\lambda(X)\geq 2\sqrt{d-1}$ by the Alon-Boppana bound (e.g. [5]), and even if $d$ grows as $N\to\infty$ (say, $\omega(1)=d<\frac{N}{2}$), one can prove that $\lambda(X)=\Omega(\sqrt{d})$ as $N\to\infty$. Hence, Ramanujan graphs realize the optimally small non-trivial eigenvalues in absolute value (up to constant). It is known that (almost) Ramanujan graphs enjoy very rich properties such as high global connectivity, pseudo-randomness, and rapid mixing. An abelian Ramanujan graph is a Cayley graph over an abelian group which is a Ramanujan graph. For a more detailed exposition of Cayley and Ramanujan graphs, see, for instance, [5].
@@ 4. A new explicit construction of $2$-quasi-perfect Lee codes
To illustrate the distinct algebraic structures of our new codes, we exhibit their parity-check matrices alongside those of the Mesnager-Tang-Qi codes for several small parameters.
@@
For $p=19\equiv 7\pmod{12}$ and $k=1$, the parity check matrix of the code $\mathcal{C}(\mathbb{F}_{19}^{2},H_{19})$ is
@@
Note that $a=0$ if and only if y=xy=-x, equivalent to $b=0$, implying that there are exactly $\frac{q-1}{2}$ unordered pair $\{x,y\}$ satisfying (4.1) with $(a,b)=(0,0)$. If $a\neq 0$, then $(a,b)$ corresponds to an unordered pair $\{x,y\}$ satisfying (4.1) and $x+y\neq 0$ by a one-to-one mapping. Indeed, by (4.1), we have $xy=\frac{a^{3}-b}{3a}$, and hence $x,y$ must be roots of the quadratic equation $t^{2}-at+\frac{a^{3}-b}{3a}=0$. The total number of unordered pairs $\{x,y\}$ from $\mathbb{F}_{q}^{*}$ (allowing $x=y$) is equivalent to choosing 2 elements from q1q-1 elements with replacement, which is $\binom{(q-1)+2-1}{2}=\binom{q}{2}=\frac{q(q-1)}{2}$. Thus, the number of unordered pairs with $x+y\neq 0$ is:
$\frac{q(q-1)}{2}-\frac{q-1}{2}=\frac{(q-1)^{2}}{2}.$
Hence, we conclude that $H_{q}^{(2)}=\frac{(q-1)^{2}}{2}+1$ by adding $(0,0)$. The proof completes after showing that $H_{q}\cap H_{q}^{(2)}=\emptyset$ as $H_{q}^{(0)}\subset H_{q}^{(2)}$ and $H_{q}^{(1)}=\#H_{q}=q-1$. Suppose that there exists $(a,a^{3})\in H_{q}$ such that there are $x,y\in\mathbb{F}_{q}^{*}$ with
@@
This implies that $3axy=0$, which is impossible as $x,y,a\neq 0$ and $p\geq 5$. ∎
@@
As the statement for the case $b=a^{3}$ is trivial, suppose $b\neq a^{3}$ and $F_{a,b}$ is not absolutely irreducible. Then, $F_{a,b}$ must have a linear factor which is either x+ycx+y-c, xcx-c or ycy-c for some constant $c$. If x+ycx+y-c is a factor of $F_{a,b}$, then substituting $x+y=c$ into $F_{a,b}(x,y)=0$ yields that $cxy-ac^{2}+a^{2}c+\frac{b-a^{3}}{3}=0$ holds for any $x$ and $y$. Hence we have $b=a^{3}$, a contradiction. If xcx-c is a factor of $F_{a,b}$, then it holds for any $y$ that $cy(c+y)-a(c+y)^{2}+a^{2}(c+y)+\frac{b-a^{3}}{3}=0$. Hence we have $c=a$ and $b=a^{3}$, which cannot occur by the assumption. By the same discussion, ycy-c cannot be a factor of $F_{a,b}$, completing the proof. ∎
@@
If $b=a^{3}$, as shown in Lemma 12, the curve $C_{a,b}$ contains the line $x+y=0$. Since $a\neq 0$ holds by the assumption, all points (t,t)(t,-t), $t\in\mathbb{F}_{q}^{*}$, on $C_{a,b}$ are desired ones.
@@ 2. $(2)$
$Q(x,y)=\left(x+\frac{y}{2}\right)^{2}+\frac{3}{4}y^{2}.$
Let $u=x+\frac{y}{2}$ and $v=y$. Then $Q(x,y)$ is equivalent to the diagonal form $u^{2}+\frac{3}{4}v^{2}$.
  • If 3-3 is a square in $\mathbb{F}_{q}$, then $\frac{3}{4}$ can be written as $s^{2}$ for some $s\in\mathbb{F}_{q}^{*}$. Thus, $Q$ is equivalent to $u^{2}-s^{2}v^{2}=(u-sv)(u+sv)$, which is equivalent to $Q_{-}(x,y)=xy$. In this case, $\mathcal{C}(\mathbb{F}_{q}^{2},E_{Q})$ coincides with $\mathcal{C}(\mathbb{F}_{q}^{2},E_{Q_{-}})$ up to coordinate permutations and sign changes.
  • If 3-3 is a non-square in $\mathbb{F}_{q}$, then $\frac{3}{4}$ is a non-square (since $\frac{1}{4}$ is a square). Let $\delta=-\frac{3}{4}$. Then $Q$ is equivalent to $u^{2}-\delta v^{2}$, which is exactly $Q_{+}(u,v)$. In this case, $\mathcal{C}(\mathbb{F}_{q}^{2},E_{Q})$ coincides with $\mathcal{C}(\mathbb{F}_{q}^{2},E_{Q_{+}})$ up to coordinate permutations and sign changes.
6. Concluding remarks
  • It would be of significant interest to investigate whether other abelian (almost) Ramanujan graphs harbor generating sets that yield new families of quasi-perfect Lee codes. Such discoveries would further solidify the profound connection between the theory of Lee codes and spectral graph theory.
  • The 2-quasi-perfect Lee codes presented in this paper, as well as those in [15], are derived from abelian (almost) Ramanujan graphs that are spectrally indistinguishable from random graphs of the same edge density. That is, their normalized eigenvalue distributions converge to the semicircle distribution, a hallmark of typical random graphs. In this sense, these graphs exhibit strong pseudo-randomness. Conversely, the property of 2-quasi-perfectness imposes rigid structural regularities on the underlying Cayley graph, as dictated by Lemma 2. Notably, the strictly deterministic condition in Lemma 2 (2) is not a property that random graphs possess with high probability. Therefore, we observe that spectral indistinguishability does not preclude the existence of the highly ordered local structures required for $2$-quasi-perfectness. This underscores a fascinating dichotomy: spectral indistinguishability alone is insufficient to guarantee “local” pseudo-randomness, revealing a structural gap akin to the phenomena discussed in [7].